# Generative Photography: Scene-Consistent Camera Control for Realistic Text-to-Image Synthesis

### v2

@@ Abstract
Image generation today can produce somewhat realistic images from text prompts. However, if one asks the generator to synthesize a particular camera setting such as creating different fields of view using a 24mm lens versus a 70mm lens, the generator will not be able to interpret and generate scene-consistent images. This limitation not only hinders the adoption of generative tools in photography applications but also exemplifies a broader issue of bridging the gap between the data-driven models and the physical world. In this paper, we introduce the concept of Generative Photography, a framework designed to control camera intrinsic settings during content generation. The core innovation of this work are the concepts of Dimensionality Lifting and Contrastive Camera Learning, which achieve continuous and consistent transitions for different camera settings. Experimental results show that our method produces significantly more scene-consistent photorealistic images than state-of-the-art models such as Stable Diffusion 3 and FLUX.
Figure 1: This paper introduces generative photography, a new paradigm for text-to-image generation that maintains a consistent base scene while modifying only the camera settings to achieve varied photographic effects. Current state-of-the-art text-to-image generation models like Stable Diffusion 3 (SD3) [4] and FLUX [3] face two major limitations: the inability to accurately interpret camera-specific settings and the challenge of maintaining consistency in the base scene. This paper introduces a novel approach that addresses these issues, enabling precise camera setting control and consistent scenes in generative models.
22footnotetext: This work is supported, in part, by the National Science Foundation under grants 2133032, 2134209, 2030570, and 2431505.33footnotetext: Project page: https://generative-photography.github.io/project/.
@@ 1 Introduction
Since the breakthrough of probabilistic diffusion models in the early 2020’s [23, 59, 51, 15], foundational vision models have created unprecedented opportunities for artificially generated content [5, 7, 72, 20, 33, 73, 52, 17, 51, 49, 53, 43, 4, 3]. Content generation processes that would have taken days of human labor in the past are now possible by machines in just a few seconds. However, while today’s text-to-image engines can generate somewhat realistic images, professional photographers still have many reservations about these tools because even the most basic camera effects cannot be produced. For example, the prompt “quiet mountain trail, 24mm lens” and another prompt “quiet mountain trail, 70mm lens” will have no differences in the field of view (FoV) but different rocks and trees, as illustrated in Fig. 1 (b).
@@
The inability of current diffusion models to effectively comprehend camera settings is not merely a matter of scaling laws[29], but rather a fundamental disconnect between data-driven models and the physical world. Recent research [28] demonstrates that increasing training data and model parameters are insufficient for generation models to capture essential physics principles. To bridge this gap, we need to “teach physics” to the models. However, teaching physics is not a trivial task because we lack both training data and suitable mathematical tools to disentangle camera physics from the scene content. The goal of this paper is to propose a new framework, driven by a co-design of data and network architecture, to enable generative photography, specifically focusing on achieving image generation with intrinsic camera setting awareness.
What is Generative Photography? Generative photography is a new paradigm in photography where content is generated instead of captured. Generative photography stresses camera-awareness. On top of an existing text-to-image generation process, we demand the model to comprehend the typical camera settings: adjusting the aperture, shutter speed, focal length, and color temperature. By adjusting these camera settings, we can create a variety of photorealistic effects such as different bokeh effects, exposure, color, and zoom. A successful generative photography method should satisfy three objectives: (1) the camera effects are realistically rendered; (2) by changing the camera settings, the content of the scene is not altered, e.g., the buildings remain the same buildings and persons remain the same persons; (3) adding camera awareness does not degrade the image quality when compared to the baseline models that do not possess this property.
Why do Existing Models Fail? There are two primary reasons why existing generative models fail.
  1. 1.
@@
    The available training data is very limited. To train a generative photography model, we need sets of images captured under different camera settings, such as varying apertures or focal lengths. Apart from the fact that acquiring these paired images is tedious and time/manpower-consuming, there is also a lack of textual labels. Therefore, in the absence of sufficient training data, it is challenging to train or fine-tune existing visual models.
  2. 2.
    Even if we have the training data, it remains unclear how one can disentangle the information of scene embeddings from the camera embeddings. If we cannot disentangle these embeddings, then it will be impossible to keep the scene unchanged while altering the camera settings. This would lead to severe scene consistency problems as illustrated in Fig. 1.
In this paper, we emphasize the concept of scene consistency from one camera setting to another camera setting. For example, if we change from “24mm lens” to “70mm lens”, there is a smooth transition from one field of view (FoV) to another. Such a smooth change creates a camera dimension, similar to the concept of spatial dimension and temporal dimension that we used to see. Therefore, to process this camera dimension, we should leverage two approaches:
  • Dimensionality Lifting. We should lift the space-only problem to a space-camera joint problem. The lifting and the joint optimization will enable new disentanglement capabilities and ensure consistency.
  • Contrastive Camera Learning. We should design a contrastive camera learning strategy to facilitate the generative model’s understanding of camera physics. This strategy operates at both the data and network levels, jointly emphasizing the changes introduced by varying camera settings within the same scene.
@@
    To support generative photography, we present two new techniques, dimensionality lifting and contrastive camera learning, to allow scene-consistent camera control for realistic text-to-image generation. This framework can be used on various camera effects, producing appealing camera control with consistent scenes.
@@ 2.2 Camera-Awareness
  • Camera Intrinsic Settings. When it comes to intrinsic camera settings, there is very limited work except for some basic zoom-in/out effects [20, 7, 2, 60]. One of the main challenges is that most datasets used for training vision models often lack comprehensive camera settings. Even when some camera metadata is included, there is typically an absence of multi-setting metadata for the same scene. This lack of parameter variation means that the available metadata cannot fully capture the true physical meaning of these settings. For instance, for a given scene, having only a single 50mm f/8 setting makes it difficult to infer how the depth of field (DoF) and field of view (FoV) would change under a 16mm f/4 setting for the same scene.
@@ 2.3 Information Disentanglement
Beyond the previous related works, a major difficulty in camera-aware content generation is the user-computer interface. In today’s text-to-image generation, a user needs to modify the prompt to control the image attributes. However, once the camera prompt is modified, the overall image structure can be severely distorted due to the sensitivity of outputs to the prompt-seed combination [31, 52, 66, 18]. This happens because we are not able to disentangle the camera embeddings and scece embeddings.
@@
Figure 2: (a) Existing T2I models struggle to perceive physical camera settings and maintain consistency across multiple settings, even when the random seed is fixed. (b) We solve this problem by lifting camera-controlled text-to-image (T2I) generation into text-to-video (T2V) generation, thereby decoupling scene description from camera settings and achieving better scene consistency.
In this paper, we introduce the task of generative photography. We propose a framework consisting of two concepts: dimensionality lifting and contrastive camera learning, which disentangle camera and scene embeddings. Our framework enables precise control over various camera settings, such as shutter speed, aperture, focal length, and color temperature, while maintaining scene consistency, allowing for exposure, bokeh rendering, zoom, and color control in generated images. These aspects will be detailed in the proposed methods section.
@@ 3.1 Dimensionality Lifting
3.2 Contrastive Camera Learning
@@
Another challenge in generative photography is how to teach the model to understand the physical settings of a camera. To achieve this, we propose a contrastive camera learning approach, consisting of two primary components: contrastive data and contrastive camera encoder. On the data side, we build a contrastive dataset that provides image-text pairs of a fixed scene under varying values of specific camera settings. At the network level, we inject the differences between adjacent camera settings to help the network further understand the impact of the variations.
3.2.1 Contrastive Data
@@
The contrastive dataset provides data-driven support for camera settings understanding in generative tasks.
@@
What is Contrastive Data and Why is it Useful? Contrastive data involves creating pairs or sets of images that highlight specific differences, such as varying camera settings, while keeping other elements constant. Contrastive data is advantageous as it enables models to focus on and learn from the differences between images, rather than being overwhelmed by the vast diversity of real-world data. Here is a simple example of contrastive data.
@@
Example. If the images vary by color temperature, their corresponding textual labels might be: Invariant scene description: "A squirrel eating a leaf from a tree." Color temperature for each frame: <9933K; 3626K; 6302K; 4039K; 2400K >
@@
How is the Contrastive Data Constructed? We propose to construct a contrastive dataset in three steps following the outline shown in Fig. 3.
Figure 3: The pipeline of building contrastive data.
@@
(i). Vision Language Model Captioning for Base Images: To ensure that the data we generate can be used for training, we first need to collect base images and obtain their textual labels. The base images we select must be of high quality and meet specific requirements for camera generation tasks. For example, for shutter speed control, the base images should have a wide dynamic range and appropriate exposure; for color temperature control, all base images should lean towards a neutral color temperature; for bokeh rendering tasks, the base images should be all-in-focus images; and for focal length control, the base images need to have sufficiently high resolution. We employ a vision language model (VLM) LLaVA [37] to generate scene captions for a base image, which we use as the invariant scene descriptions for this contrastive set.
@@
(ii). Random Sampling on Continuous Camera Space: Camera settings are typically treated as continuous variables, such as focal length, shutter speed, and color temperature. We argue that each value on the camera setting scale is crucial, as it embodies a deeper physical understanding rather than merely representing a few discrete values. Continuous sampling supports continuous-scale training [8, 10, 64], which enhances the model’s ability to perform across any value on the scale. Therefore, when constructing contrastive data, we perform random sampling across the continuous camera setting scale to gather multiple values. Specifically, as shown in Fig. 3, we assume an invariant scene description on a continuous camera setting scale (here exemplified by color temperature, with the vertical axis representing color temperature). For each instance, we sample several (5 here) points along the horizontal axis and use this collection of points as a set to describe camera setting variations. We did not order the randomly collected data points in each set, ensuring that our network does not merely fit a specific directional pattern (such as simple zoom in/out for focal length) but instead learns the numerical values themselves. For the camera’s color temperature, we collect a range from 2,000 to 10,000 Kelvin. For shutter speed, we re-normalize it to a scale of 0.1 to 1.0 (where larger values correspond to brighter images). For focal length, we sample values between 24 mm and 70 mm. Regarding bokeh rendering, we sample blur parameter values from 1 to 30, with larger values corresponding to stronger blur effects.
@@
3.2.2 Contrastive Camera Encoder
Figure 4: The overall architecture of contrastive camera encoder.
As shown in Fig 2 (b), after dimensionality lifting, we use the scene description as the primary input to the T2V model to represent the base scene, with the camera settings as additional conditions. These settings are fed into the Contrastive Camera Encoder, which injects them into the foundation models to control the camera effects for each frame. In the contrastive camera encoder, we enhance the network’s ability to understand and generate camera settings by embedding camera settings and providing additional differential information.
@@ 3.2.2 Contrastive Camera Encoder
We propose that the Contrastive Camera Encoder enhances the model’s ability to differentiate camera settings by focusing on comparative effects between them. This approach moves beyond a solely data-driven paradigm, significantly reducing the training data requirements and computational costs, while providing the network with more robust and precise control capabilities.
@@ 4.1 Implementation Details
Training and Inference. In this work, we select AnimateDiff [20] as our T2V base model, reducing the generated video frames to 5 to decrease computational costs for both training and inference. During training, most parameters of AnimateDiff are kept frozen, with only the motion LoRA [26] and our contrastive camera encoder being fine-tuned/trained. We use a learning rate of 1e-4 with the Adam optimizer. The video resolution for training is set to 256 × 384. For each task, we train with a batch size of 8 for 25,000 epochs, taking approximately 10 hours on an Nvidia A100 80GB GPU.
@@
Metrics. Our evaluation metrics primarily focus on accuracy to physical laws and scene consistency. The accuracy of generated images with respect to camera physical settings is evaluated using the Pearson correlation coefficient (CorrCoef) of trend changes between the generated images and the reference video. For scene consistency, we use frame-wise Learned Perceptual Image Patch Similarity (LPIPS) [74] to calculate the perceptual feature distance between frames generated with different camera settings. Additionally, to evaluate the impact of added camera controls on generation quality, we use CLIP [48] to compute the similarity score between generated images and the text prompt. For more details on the metrics, please refer to the supplementary material.
@@ 4.2 Comparisons with Other Methods
Accuracy Consistency Quality Accuracy Consistency Quality Accuracy Consistency Quality Accuracy Consistency Quality
@@
Table 1: Quantitative comparison with different generative methods. Accuracy is computed by comparing with reference videos; Consistency by frame-wise LPIPS; Quality by CLIP. Reference refers to the results obtained from physical simulations. FT denotes fine-tuning on our contrastive data. We highlight the best and second-best values for each metric
@@
Accuracy Consistency Quality Accuracy Consistency Quality Accuracy Consistency Quality Accuracy Consistency Quality
@@
To evaluate our method, we conduct comparisons against three categories of approaches: first, the state-of-the-art text-to-image generation models, Stable Diffusion 3 (SD3) [4] and FLUX [3]; second, the text-to-video generation model, AnimateDiff [20], fine-tuned on our contrastive dataset to assess the performance of a purely data-driven approach; and finally, CameraCtrl [22], where we adapt the model to control intrinsic camera settings instead of external ones, and train it on our dataset. Together, these baseline models form a comprehensive evaluation framework, encompassing image and video generation, external camera parameter control, and other related aspects.
In Fig. 5 demonstrates that our method generates images with excellent scene consistency, while realistically simulating camera effects based on any given setting value. Notably, in bokeh rendering, although no depth information was provided during inference, the rendered scenes appear to exhibit depth awareness, consistently keeping the foreground sharp while only the background varies according to the bokeh blur parameter. In color temperature control, our model also shows fine-grained control capabilities, producing images at 3, [^0] K and 3,002K with only minimal differences in color temperature. From Table 1, although SD3 and FLUX demonstrate a clear advantage in generation quality, it exhibits lower consistency and accuracy between frames with different camera settings. After fine-tuning using our comparative dataset, both AnimateDiff and CameraCtrl showed obvious improvements in accuracy. Our method shows a significant advantage in both the accuracy and consistency of generated camera settings. It also performs well in generation quality, suggesting that our approach maintains the model’s overall generative capabilities without significant compromise. Here, a lower LPIPS score is not always better, as variations in color temperature, exposure, focal length, and bokeh can affect frame differences. The key metric is proximity to the reference videos.

### v3

@@ Abstract
Image generation today can produce somewhat realistic images from text prompts. However, if one asks the generator to synthesize a specific camera setting such as creating different fields of view using a 24mm lens versus a 70mm lens, the generator will not be able to interpret and generate scene-consistent images. This limitation not only hinders the adoption of generative tools in professional photography but also highlights the broader challenge of aligning data-driven models with real-world physical settings. In this paper, we introduce Generative Photography, a framework that allows controlling camera intrinsic settings during content generation. The core innovation of this work are the concepts of Dimensionality Lifting and Differential Camera Intrinsics Learning, enabling smooth and consistent transitions across different camera settings. Experimental results show that our method produces significantly more scene-consistent photorealistic images than state-of-the-art models such as Stable Diffusion 3 and FLUX. Our code and additional results are available at https://generative-photography.github.io/project.
Figure 1: This paper introduces generative photography, a new paradigm for text-to-image generation that maintains a consistent base scene while modifying only the camera settings to achieve varied photographic effects. Current state-of-the-art text-to-image generation models like Stable Diffusion 3 (SD3) [4] and FLUX [1] face two major limitations: failure to accurately interpret camera-specific settings and difficulties in maintaining consistency in the base scene. This paper introduces a novel approach that addresses these issues, enabling precise camera setting control and maintaining scene consistency in generative models.
@@ 1 Introduction
Since the breakthrough of probabilistic diffusion models in the early 2020’s [24, 60, 52, 15], foundational vision models have created unprecedented opportunities for artificially generated content [5, 7, 73, 21, 34, 74, 53, 18, 52, 50, 54, 44, 4, 1]. Content generation processes that would have taken days of human labor in the past are now possible by machines in just a few seconds. Although current text-to-image models can produce perceptually realistic images, professional photographers remain skeptical, as these tools still cannot reliably replicate even basic camera effects. For example, the prompt “quiet mountain trail, 24mm lens” and another prompt “quiet mountain trail, 70mm lens” will have no differences in the field of view (FoV) but different rocks and trees, as illustrated in Fig. 1 (b).
@@
The inability of current diffusion models to accurately interpret camera settings reflects not merely limitations of scaling laws [30], but a fundamental gap between data-driven models and physical-world principles. Recent research [29] demonstrates that increasing training data and model parameters are insufficient for generation models to capture essential physics principles. To bridge this gap in photography context, we aim to teach camera physics to the models. However, teaching camera physics is not a trivial task due to the lack of training data and the fundamental problem of existing image/video generation foundation models [29]. This paper introduces a novel framework, jointly designed through data curation and network architecture, to enable Generative Photography—specifically targeting image generation with camera intrinsics controls.
@@
Why do Existing Generative Models Fail to Understand Camera Physics? There are two primary reasons why existing generative models fail: (1) The available training data is very limited. To train a generative photography model, we need sets of images captured under different camera settings, such as varying apertures or focal lengths. Apart from the fact that acquiring these paired images is tedious and time/manpower-consuming, there is also a lack of textual labels. Therefore, in the absence of sufficient training data, it is challenging to train or fine-tune existing visual models; (2) Even if we have the training data, it remains unclear how one can disentangle the information of scene embeddings from the camera embeddings. If we cannot disentangle these embeddings, then it will be impossible to keep the scene unchanged while altering the camera settings. This would lead to severe scene consistency problems as illustrated in Fig. 1.
What is Generative Photography? Generative photography is an emerging approach in photography where content is generated instead of physically captured. Generative photography stresses camera-awareness. On top of an existing text-to-image generation process, we demand the model to comprehend the typical camera settings: adjusting the aperture, shutter speed, focal length, and color temperature. By adjusting these camera settings, we can generate a variety of photorealistic effects such as different bokeh effects, exposure, color, and zoom, unlocking new applications in photography. A successful generative photography method should satisfy three objectives: (1) the camera effects are realistically rendered; (2) by changing the camera settings, the content of the scene is not altered, e.g., the buildings remain the same buildings and persons remain the same persons; (3) adding camera awareness does not degrade the image quality when compared to the baseline models that do not possess this property.
In this paper, we highlight the importance of maintaining scene consistency across varying camera settings. For example, transitioning from a "24mm lens" to a "70mm lens" should result in a smooth and consistent change in the field of view (FoV). This seamless transition between different camera intrinsics introduces the concept of a camera dimension. Analogous to spatial and temporal dimensions, generative photography is to solve consistency problem in spatial and camera dimensions. To handle this camera dimension effectively, we propose two complementary approaches: Dimensionality Lifting, transforming text-to-image generation from a purely spatial problem into a joint space-camera domain for improved camera intrinsics disentanglement and spatial consistency; and Differential Camera Intrinsics Learning, a strategy designed to explicitly capture differences in camera intrinsic settings, operating simultaneously at both data and network architecture levels to reinforce consistent scene representation.
@@
    To support generative photography, we present two new techniques, dimensionality lifting and differential camera intrinsics learning, to allow scene-consistent camera control for realistic text-to-image generation. This framework can be used on various camera effects, producing appealing camera control with consistent scenes.
@@ 2.2 Camera-Awareness
Camera Intrinsic Settings. When it comes to intrinsic camera settings, there is very limited work except for some basic zoom-in/out effects [21, 7, 3, 61, 17]. One of the main challenges is that most datasets used for training vision models often lack comprehensive camera settings. Even when some camera metadata is included, there is typically an absence of multi-setting metadata for the same scene [17]. This lack of parameter variation means that the available metadata cannot fully capture the true physical meaning of these settings. For instance, for a given scene, having only a single 50mm f/8 setting makes it difficult to infer how the depth of field (DoF) and field of view (FoV) would change under a 16mm f/4 setting for the same scene.
@@ 2.3 Information Disentanglement
Beyond the previous related works, a major difficulty in camera-aware content generation is the user-computer interface. In today’s text-to-image generation, a user needs to modify the prompt to control the image attributes. However, once the camera prompt is modified, the overall image structure can be severely distorted due to the sensitivity of outputs to the prompt-seed combination [32, 53, 67, 19]. This happens because we are not able to disentangle the camera embeddings and scene embeddings.
@@
Figure 2: (a) Existing text-to-image (T2I) models struggle to perceive physical camera settings and maintain consistency across multiple settings, even when the random seed is fixed. (b) We solve this problem by lifting camera-controlled text-to-image (T2I) generation into text-to-video (T2V) generation, thereby decoupling scene description from camera settings and achieving better scene consistency.
In this paper, we introduce the task of generative photography. We propose a framework consisting of two concepts: dimensionality lifting and differential camera intrinsics learning, which disentangle camera and scene embeddings. Our framework enables precise control over various camera settings, such as shutter speed, aperture, focal length, and color temperature, while maintaining scene consistency, allowing for exposure, bokeh rendering, zoom, and color control in generated images. These aspects will be detailed in the proposed methods section.
@@ 3.1 Dimensionality Lifting
3.2 Differential Camera Intrinsics Learning
@@
Another challenge in generative photography is teaching the model to comprehend the physical settings of a camera. To address this, we introduce a differential camera intrinsics learning approach comprising two key components: differential data and a differential camera encoder. On the data side, we construct a differential dataset containing image-text pairs of a fixed scene captured under varying camera settings. At the network level, we incorporate the differences between adjacent camera settings to enhance the model’s understanding of their effects.
3.2.1 Differential Data
@@
The differential dataset provides data-driven support for camera settings understanding in generative tasks.
@@
What is Differential Data and Why is it Useful? Differential data involves creating pairs or sets of images that highlight specific differences, such as varying camera settings, while keeping other elements constant. Differential data is advantageous as it enables models to focus on and learn from the differences between images, rather than being overwhelmed by the vast diversity of real-world data. Here is a simple example of differential data.
@@
Example. If the images vary by color temperature, their corresponding textual labels are: Invariant scene description: "A squirrel eating a leaf from a tree." Color temperature for each frame: <9933K; 3626K; 6302K; 4039K; 2400K >
@@
How is the Differential Data Constructed? We propose to construct a differential dataset in three steps following the outline shown in Fig. 3.
Figure 3: The pipeline of building differential data.
@@
(i). Vision Language Model Captioning for Base Images: To ensure that the data we generate can be used for training, we first need to collect base images and obtain their textual labels. The base images we select must be of high quality and meet specific requirements for camera generation tasks. For example, for shutter speed control, the base images should have a wide dynamic range and appropriate exposure; for color temperature control, all base images should lean towards a neutral color temperature; for bokeh rendering tasks, the base images should be all-in-focus images; and for focal length control, the base images need to have sufficiently high resolution. We employ a vision language model (VLM) LLaVA [38] to generate scene captions for a base image, which we use as the invariant scene descriptions for this differential set.
@@
(ii). Random Sampling on Continuous Camera Space: Camera settings are typically treated as continuous variables, such as focal length, shutter speed, and color temperature. We argue that each value on the camera setting scale is crucial, as it embodies a deeper physical understanding rather than merely representing a few discrete values. Continuous sampling supports continuous-scale training [8, 10, 65], which enhances the model’s ability to perform across any value on the scale. Therefore, when constructing differential data, we perform random sampling across the continuous camera setting scale to gather multiple values. Specifically, as shown in Fig. 3, we assume an invariant scene description on a continuous camera setting scale (here exemplified by color temperature, with the vertical axis representing color temperature). For each instance, we sample several (5 here) points along the horizontal axis and use this collection of points as a set to describe camera setting variations. We did not order the randomly collected data points in each set, ensuring that our network does not merely fit a specific directional pattern (such as simple zoom in/out for focal length) but instead learns the numerical values themselves. For the camera’s color temperature, we collect a range from 2,000 to 10,000 Kelvin. For shutter speed, we re-normalize it to a scale of 0.1 to 1.0 (where larger values correspond to brighter images). For focal length, we sample values between 24 mm and 70 mm. Regarding bokeh rendering, we sample blur parameter values from 1 to 30, with larger values corresponding to stronger blur effects.
@@
3.2.2 Differential Camera Encoder
Figure 4: The overall architecture of differential camera encoder.
As shown in Fig 2 (b), after dimensionality lifting, we use the scene description as the primary input to the T2V model to represent the base scene, with the camera settings as additional conditions. These settings are fed into the Differential Camera Encoder, which injects them into the foundation models to control the camera effects for each frame. In the differential camera encoder, we enhance the network’s ability to understand and generate camera settings by embedding camera settings and providing additional differential information.
@@ 3.2.2 Differential Camera Encoder
We propose that the Differential Camera Encoder enhances the model’s ability to differentiate camera settings by focusing on comparative effects between them. This approach moves beyond a solely data-driven paradigm, significantly reducing the training data requirements and computational costs, while providing the network with more robust and precise control capabilities.
@@ 4.1 Implementation Details
Training and Inference. In this work, we select AnimateDiff [21] as our T2V base model, reducing the generated video frames to 5 to decrease computational costs for both training and inference. During training, most parameters of AnimateDiff are kept frozen, with only the motion LoRA [27] and our differential camera encoder being fine-tuned/trained. We use a learning rate of 1e-4 with the Adam optimizer. The video resolution for training is set to 256 × 384. For each task, we train with a batch size of 8 for 25,000 epochs, taking approximately 10 hours on an Nvidia A100 80GB GPU.
@@
Metrics. Our proposed evaluation metrics primarily focus on accuracy to physical laws and scene consistency. The accuracy of generated images with respect to camera physical settings is evaluated using the Pearson correlation coefficient (CorrCoef) of trend changes between the generated images and the reference video. For scene consistency, we use frame-wise Learned Perceptual Image Patch Similarity (LPIPS) [75] to calculate the perceptual feature distance between frames generated with different camera settings. Additionally, to evaluate the impact of added camera controls on generation following, we use CLIP [49] to compute the similarity score between generated images and the text prompt. For more details on the metrics, please refer to the supplementary material.
@@ 4.2 Comparisons with Other Methods
Accuracy Consistency Following Accuracy Consistency Following Accuracy Consistency Following Accuracy Consistency Following
@@
Table 1: Quantitative comparison with different generative methods. Accuracy is computed by comparing with reference videos; Consistency by frame-wise LPIPS; Prompt following by CLIP. Reference refers to the results obtained from physical simulations. FT denotes fine-tuning on our differential data. We highlight the best and second-best values for each metric.
@@
Accuracy Consistency Following Accuracy Consistency Following Accuracy Consistency Following Accuracy Consistency Following
@@
To evaluate our method, we conduct comparisons against three categories of approaches: first, the state-of-the-art text-to-image generation models, Stable Diffusion 3 (SD3) [4] and FLUX [1]; second, the text-to-video generation model, AnimateDiff [21], fine-tuned on our differential dataset to assess the performance of a purely data-driven approach; and finally, CameraCtrl [23], where we adapt the model to control intrinsic camera settings instead of external ones, and train it on our dataset. Together, these baseline models form a comprehensive evaluation framework, encompassing image and video generation, external camera parameter control, and other related aspects.
Fig. 5 demonstrates that our method preserves excellent scene consistency while realistically simulating camera effects for any given setting. Notably, in bokeh rendering, although no depth information was provided during inference, the rendered scenes appear to exhibit depth awareness, consistently keeping the foreground sharp while only the background varies according to the bokeh blur parameter. In color temperature control, our model also shows fine-grained control capabilities, producing images at 3, [^0] K and 3,002K with only minimal differences in color temperature. From Table 1, although SD3 and FLUX demonstrate a clear advantage in prompt following, it exhibits lower consistency and accuracy between frames with different camera settings. After fine-tuning using our comparative dataset, both AnimateDiff and CameraCtrl showed obvious improvements in accuracy. Our method shows a significant advantage in both the accuracy and consistency of generated camera settings. It also performs well in prompt following, suggesting that our approach maintains the model’s overall generative capabilities without significant compromise. Here, a lower LPIPS score is not always better, as variations in color temperature, exposure, focal length, and bokeh can affect frame differences. The key metric is proximity to the reference videos. Compared to other base models, our approach still excels in prompt following, indicating that the additional camera control does not compromise generative performance.